# Enhancing Hemodynamic Parameter Estimations: Nonlinear Blood Behavior in 4D Flow MRI

### v2

@@ Abstract
Hemodynamic parameters are often estimated assuming a constant Newtonian viscosity, even though blood exhibits shear-thinning behavior. This article investigates the influence of blood rheology and hematocrit (Hct) percentage on the estimation of Wall Shear Stress (WSS), rate of viscous Energy Loss ($\dot{E}_{L}$) at different points in the cardiac cycle, and the Oscillatory Shear Index (OSI). We focus on a hematocrit-dependent power-law non-Newtonian model, considering a wide range of Hct values at physiological temperature, with rheological parameters obtained from previously reported experimental data. In all cases, we systematically compared WSS, $\dot{E}_{L}$, and OSI using both Newtonian and power-law models, underscoring the crucial role of blood rheology in accurately assessing cardiovascular diseases. Our results show that, in in-silico experiments, differences in WSS and $\dot{E}_{L}$ across a wide range of Hct values can reach as high as 190% and 113% at systole, and as low as -72% and -74% at diastole, respectively. In in-vivo data, differences in WSS and $\dot{E}_{L}$ can reach up to -45% and -60% at systole, and range from -69% to 73% at diastole. This study enhances our understanding of the impact of blood rheology on hemodynamic parameter estimations using both in-silico and in-vivo aortic 4D Flow MRI data.
keywords:
4D Flow MRI , Blood Rheology , Cardiovascular MRI , Hemodynamics , Finite Elements
journal: Biomedical Signal Processing and Control
[…]
[EIEPUCV]organization=School of Electrical Engineering, Pontificia Universidad Católica de Valparaíso,addressline=Av. Brasil 2147, city=Valparaíso, postcode=2340000, country=Chile
[…]
[EICPUCV]organization=School of Civil Engineering, Pontificia Universidad Católica de Valparaíso,addressline=Av. Brasil 2147, city=Valparaíso, postcode=2340000, country=Chile
[…]
[JAPAN]organization=Department of Radiology, Nippon Medical School Musashi Kosugi Hospital,city=Kanagawa, postcode=2118533, country=Japan
[…]
[UTFSM]organization=Departamento de Informática, Universidad Técnica Federico Santa María,city=Santiago, postcode=8940000, country=Chile
[…]
[USACH]organization=Department of Mechanical Engineering, Universidad de Santiago de Chile,city=Santiago, postcode=9160000, country=Chile
@@ 1 Introduction
Hemodynamic parameters estimated from 4D Flow Magnetic Resonance (MR) images such as Wall Shear Stresses (WSS), Oscillatory Shear Indexes (OSI), and viscous Energy Loss ($E_{L}$) have been successfully used for the assessment of several cardiovascular diseases in large vessels, including Bicuspid Aortic Valve [1], aortic coarctations [2], stenosis [3], among others [4]. These parameters need as input a blood viscosity $\mu$, which in most of the cases is assumed constant and within the range of $0.03$ to $0.045$ poise [5, 6, 7, 8, 9, 10]. This assumption has been accepted in the cardiovascular MRI community for the past decades, serving as the gold standard for the estimation of viscosity-dependent hemodynamic parameters from 4D Flow MR images. However, the blood viscosity has an intrinsic shear-thinning nonlinear behavior (i.e., viscosity decreases as the shear rates increases) that, to the author’s knowledge, has not been systematically considered in estimating these parameters.
Blood is a non-homogeneous and non-linear fluid comprising red blood cells, white blood cells, and platelets suspended in plasma. The behavior of blood can be modeled using continuum mechanics and experimental data, leading to non-Newtonian constitutive models. Generally, blood exhibits shear-thinning behavior, primarily dependent on the hematocrit (Hct) percentage. This shear-thinning behavior can be effectively described using models such as the power-law [11, 12], Carreau-like [13, 14, 15], Cross [16, 12], and Casson models [17]. Considering that the power-law has been successfully fitted to a wide range of Hct percentages under physiological temperatures [18, 19, 20], we have adopted this model as a suitable candidate for explaining blood’s fluid behavior.
Increased Hct can result from dehydration and polycythemia, while decreased values can be caused by anemia, overhydration , kidney failure, pregnancy, or chronic inflammatory conditions [21, 18, 19, 20, 22]. Although Hct can range from $10\%$ to $70\%$ in individuals with these conditions, the normal range is $47\pm 7\%$ for men and $42\pm 6\%$ for women [23]. Besides, shear rate depends on spatial and temporal blood flow characteristics, opening up the question of whether Newtonian modeling, accepted by the MRI community, truly represents a whole cardiac cycle. In this regard, including a non-Newtonian rheological behavior could generate potential differences in hemodynamic parameters across patients, especially in non-physiological blood conditions.
The influence of blood rheology has been studied in various hemodynamic contexts using 4D Flow MRI. For example, a shear-thinning viscosity model dependent on Hct was employed to estimate hemodynamic parameters from in-vivo 4D Flow data in the circle of Willis in children and adults with sickle cell disease [24], as well as for flow and hemodynamic analysis of the left ventricle [25]. In the latter study, two different rheology models were assessed and compared to state-of-the-art Newtonian viscosity estimations, revealing notable differences between models. In-silico studies, involving geometries resembling Fontan patients, as well as in-vitro 4D Flow MRI phantom experiments, were conducted in [26] to evaluate the impact of shear-dependent fluid viscosity on in-vitro Fontan circulation. A similar study assessed prosthetic heart valves using in-vitro experimental data obtained through Particle Imaging Velocimetry [27].
From a numerical standpoint, the hemodynamics community has largely advanced using the finite element method (FEM) [28, 29]. FEM has been extensively applied to analyze and solve the governing dynamics of the cardiovascular system, such as the Navier-Stokes equations and related models [30], and to compute full states and related parameters for inverse problems based on medical data [31, 32]. In [33], a numerical study investigated the influence of hematocrit levels in blood, considering both anemic and physiological patients in a simplified carotid artery model. The study found differences in vortex shedding frequency and flow bifurcation between the carotid outlet branches. In [34], a numerical analysis of aortic aneurysms evaluated non-Newtonian and Newtonian models through CFD simulations with two turbulence models, revealing small differences in flow dynamics and hemodynamic parameters across viscosity and turbulence models. Similarly, [35] assessed the arterial and venous systems through numerical experiments with different viscosity models, concluding that accounting for non-Newtonian behavior is crucial for cardiovascular flows exhibiting low shear rates and recirculation.
@@
Despite the valuable insights provided by the aforementioned hemodynamics works where the shear-thinning rheological behavior has been incorporated in the blood flow modeling, it is essential to mention some limitations and open questions. The effect of rheological behavior in large vessels like the aorta has not been analyzed in detail [24, 25]. The impact of blood hematocrit has not been studied in wide ranges [26, 35, 34], or has not addressed their influence on specific hemodynamic parameters like WSS and OSI [27]. Additionally, none of these studies have evaluated whether, under common MR imaging conditions and issues such as partial volume effects, flow artifacts, and noise, the impact of the shear-thinning effect is still measurable.
In this work, we demonstrate that if blood exhibits shear-thinning behavior, this effect can be accurately measured from 4D Flow MRI data, even in the presence of imaging artifacts under common conditions. To support this claim, we conducted numerical and in-vivo experiments where hemodynamic parameters were estimated from 4D Flow data using both power-law and linear viscosity models for blood rheology. Comparisons were made for WSS, $\dot{E}_{L}$ (rate of viscous $E_{L}$), and OSI estimations. Building upon these state-of-the-art models, this study integrates the 3D Navier-Stokes equations with 0D surrogate models for downstream flow and non-linear stress tensors to simulate shear-thinning blood behavior in experimental settings. Additionally, in-vivo experiments incorporate power-law viscosities adjusted to each patient’s Hct, measured at physiological temperatures. This work contributes to a deeper understanding of blood rheology in MRI and CFD studies, laying the groundwork for more specific future investigations.
@@ 2.1 Mathematical modeling
Let $\Omega\in\mathbb{R}^{3}$ be the three-dimensional computational domain, whose two-dimensional boundaries are decomposed in disjoint subsets such that $\partial\Omega=\Gamma_{\text{in}}\cup_{k=1}^{l}\Gamma_{\text{out}}^{k}\cup\Gamma_{\text{wall}}$, referring $\Gamma_{\text{out}}$ for the outlets, $\Gamma_{\text{in}}$ for the inlet, and $\Gamma_{\text{wall}}$ for the artery walls. The incompressible Navier-Stokes (NS) equations relates the velocity vector field ($\boldsymbol{u}$) and the pressure ($p$) field as follows:
@@
where $k=1,\ldots,l$ is the outlet counter, $\tau$ denotes the shear stress tensor of the fluid, and $\rho$ is the fluid density. We will discuss about modeling choices for $\tau$ in the following section. Additionally, $g(x,t)$ is an arbitrary spatially and temporarily dependent boundary condition used to define the pulsatile behavior of blood flow. To model the downstream fluid we couple the 3D-NS equations with the reduced 0D three-parameters Windkessel models [28] (see Figure 1), as it is typically done in the literature [36, 32, 37, 30]. This approach boils down to solving an ordinary differential equation (ODE) for each outlet domain, such that:
$p^{k}_{\text{wk}}=R_{p}Q^{k}(\boldsymbol{u})+p^{k}_{d},\quad k=1,\ldots,l,$ (2a)
$C^{k}\frac{dp^{k}_{d}}{dt}+\frac{p^{k}_{d}}{R^{k}_{d}}=Q^{k}(\boldsymbol{u}),\quad k=1,\ldots,l.$ (2b)
@@
Figure 1: Simulation workflow for generating 4D Flow MR images. First, the aortic model serves as input to the MAD library, where the Navier-Stokes equations with nonlinear viscosity are solved. Subsequently, $k$-space locations and sequence timings are generated using parameters from Table 3, and the signal equation in (4) is evaluated. Trapezoidal MRI gradients are considered for the generation of sequence timings.

### v3

# Evaluating the Impact of Blood Rheology in Hemodynamic Parameters by 4D Flow MRI in Large Vessels Considering the Hematocrit Effect

@@ Abstract
Aortic hemodynamic parameters estimated from 4D Flow Magnetic Resonance (MR) velocity measurements are often estimated using a constant Newtonian viscosity, neglecting blood’s shear-thinning behavior. The aim of this work is to estimate and assess whether Newtonian viscosity is sufficient to quantify these parameters, given the non-Newtonian nature of blood. Additionally, we demonstrate that shear-thinning effects remain observable in large vessels despite artifacts commonly present in 4D Flow MR images. To address this, we quantified the impact of blood rheology and hematocrit (Hct) on Wall Shear Stress (WSS), the rate of viscous Energy Loss ($\dot{E}_{L}$), and the Oscillatory Shear Index (OSI) based on velocity data obtained from 4D Flow MR images. Using a Hct-dependent power-law non-Newtonian model with experimentally derived rheological parameters, we analyzed these metrics across a broad range of Hct values at physiological temperatures in both in-silico and in-vivo MR datasets.
The results reveal significant differences between Newtonian and non-Newtonian models. In in-silico experiments, WSS and $\dot{E}_{L}$ differed by up to +189% and +112% at systole, with reductions of -74% and -80% at diastole, respectively, while OSI differences ranged from -23% to -30%. For in-vivo data, WSS and $\dot{E}_{L}$ deviations reached -44% and -60% at systole, ranging from -69% to +73% at diastole, with OSI differences averaging -21%. These findings highlights the importance of accounting for non-Newtonian blood rheology when estimating hemodynamic parameters from 4D Flow MR images in large vessels, enhancing the accuracy of cardiovascular disease assessments using in-vivo aortic data.
First two authors contributed equally
@@ 1 Introduction
4D (3D + time) Flow imaging is an advanced Magnetic Resonance (MR) technique that allows the measurement of spatially and temporally resolved velocity fields for any tissue [1]. During the past decades, 4D flow has been extensively used to assess a wide variety of diseases by evaluating blood flow in various organs, such as the heart [2, 3], brain [4, 5], and large vessels [6, 7].
@@
Hemodynamic parameters estimated from 4D Flow MR velocity measurements such as Wall Shear Stress (WSS), Oscillatory Shear Index (OSI), and viscous Energy Loss ($E_{L}$) have been successfully used for the evaluation of various cardiovascular diseases in large vessels, including Bicuspid Aortic Valve [6], aortic coarctations [8], stenosis [9], among others [10]. The estimation of these parameters requires the use of complex image-processing methodologies and the selection of an appropriate blood viscosity definition $\mu$, which is typically assumed to be constant, ranging between $0.03$ to $0.045$ poise [11, 12, 13, 14, 15, 16].
This assumption has been widely accepted in the cardiovascular MRI community for decades, serving as the gold standard for estimating viscosity-dependent hemodynamic parameters from velocity measurements obtained using 4D Flow MR images. However, blood viscosity has an intrinsic shear-thinning nonlinear behavior (i.e., viscosity decreases as shear rates increases), which, to the author’s knowledge, has not been systematically considered when estimating these parameters directly from blood velocity measurements obtained using 4D Flow MR imaging.
Blood is a non-homogeneous and nonlinear shear-thinning fluid composed of red blood cells, white blood cells, and platelets suspended in plasma. Its behavior can be modeled using continuum mechanics and experimental data, resulting in non-Newtonian constitutive models. This shear-thinning behavior, which depends on the shear rate, can be effectively described by models such as the power-law [17, 18], Carreau [19, 20, 21], Cross [22, 18], and Casson models [23]. The fundamental distinction between these rheological models lies in the number of constitutive parameters required to characterize the viscosity. Among these, the power-law model is the simplest one, with only two constitutive parameters required: the power-law index parameter, which defines the rheological behavior of the fluid (shear-thinning or shear-thickening), and the consistency index. The simplicity of the power-law model has been a key factor for the blood rheological community to propose empirical rheological-dependent/Hct correlations and Hct-dependent constitutive parameters for large ranges of red blood cell concentrations, a key issue for this article.
Several studies have investigated the influence of Hct on blood rheology and viscosity. For instance, in 1962, Wells and Merrill [24] examined the role of Hct in determining blood viscosity across different shear rates, demonstrating elevated viscosities at low shear rates and reduced viscosities at high shear rates. In 1981, Quemada [25] developed a viscosity model showing that increased Hct enhances Red Blood Cells (RBC) aggregation and deformation with an explicit shear-thinning behavior. In 2003, Baskurt and Meiselman [26] highlighted the exponential increase in blood viscosity with increasing Hct values due to augmented RBC aggregation, also emphasizing the blood rheology behavior. More recently, in 2020, Horner [27] provided comprehensive experimental data revealing that Hct is strongly correlated with viscoelastic and thixotropic behavior of blood, influencing rouleaux formation at low shear rates and RBC deformation at high shear rates. Furthermore, in 2021, Beris et al. [28] reinforced Hct’s central role in governing non-Newtonian behaviors, such as shear-thinning and thixotropy, through an extensive literature review.
An increase in Hct (i.e., increase in blood viscosity) can result from conditions such as dehydration and polycythemia, while a decrease (i.e., decrease in blood viscosity) can be attributed to anemia, overhydration, kidney failure, pregnancy, or chronic inflammatory diseases [29, 24, 30, 31, 32]. Although Hct can range from $10\%$ to $70\%$ in patients with these conditions, the normal range is $47\pm 7\%$ for men and $42\pm 6\%$ for women [33].
It is well known that variations in viscosity could significantly influence the interpretation of 4D Flow MRI examinations, as changes in velocity measurements might reflect altered flow dynamics driven by viscosity variations related to underlying health conditions. Moreover, the shear rate, which acts as an independent variable in all the previously discussed rheological models, is affected by the spatial and temporal characteristics of blood flow. These issues raise the question of whether the Newtonian viscosity model, which is widely accepted in the MRI community and used to estimate hemodynamic parameters directly from 4D flow measurements, is adequate for accurately characterizing the entire cardiac cycle. In this regard, the incorporation of a non-Newtonian rheological model may lead to notable differences in hemodynamic parameters estimated from 4D Flow MR images between patients, particularly under non-physiological blood composition and flow conditions. An additional consideration is whether these differences, if present, can be observed from velocity measurements obtained using 4D Flow MRI, an imaging technique that is inherently susceptible to various artifacts that can bias the estimation of these parameters [34].
Whether to use Newtonian or non-Newtonian rheological models in hemodynamics remains an open question that has been extensively explored from a numerical perspective. For instance, in [35], Apostolidis et al. investigated the effects of non-Newtonian rheology, represented by the Casson viscoplastic model, compared to its Newtonian counterpart, on a simplified left coronary artery geometry using Computational Fluid Dynamics (CFD) simulations. Both viscosity models were calibrated to an Hct of 40% (representative of a healthy subject), and their findings revealed significant differences (up to 50%) in simulation outputs, particularly in terms of WSS and peak pressure. Similarly, in [36], Elhanafy et al. studied the effects of blood shear-thinning and viscoelastic properties in an idealized abdominal aortic aneurysm using CFD. The non-Newtonian viscosity was modeled using the Carreau-Yasuda formulation, combined with the Oldroyd-B model to account for the viscoelastic component of the stress tensor, with parameters calibrated for a healthy subject. Their primary conclusion was that under low shear rate conditions, shear-thinning and viscoelastic properties of blood cannot be neglected due to the significant differences observed compared to the Newtonian model.
In [37], Fuchs et al. evaluated the significance of non-Newtonian viscosity models, including Casson, Quemada, and Walburn-Schneck, for a fixed Hct of 45% (representative of a healthy subject) in the thoracic aorta using CFD. Their findings were compared against simulations employing a fixed Newtonian viscosity, with differences assessed in parameters such as WSS, OSI, and time-averaged WSS (TAWSS). Similar to the findings of Apostolidis et al., their results demonstrated substantial regional variations in WSS between simulations using non-Newtonian and Newtonian viscosity models, with the largest differences observed in regions of low shear rate. However, for time- or spatially-averaged quantities such as OSI and TAWSS, these differences, while still noticeable, were reduced due to the filtering effect inherent in the averaging process. With the same idea of highlighting the importance of using non-Newtonian constitutive models, in [38], De Nisco et al. investigated whether shear-thinning non-Newtonian rheology is necessary for accurately modeling coronary arteries. The study compared the Carreau and Newtonian viscosity models, with parameters calibrated for an Hct of 43% (representative of a healthy subject), using CFD simulations on 144 patient-specific coronary artery geometries. The comparison focused on near-wall parameters, such as WSS, OSI, and TAWSS, as well as intravascular hemodynamic parameters, including helicity-derived metrics. Their results indicated that the assumption regarding blood rheology has a negligible impact on both WSS and helical flow profiles associated with coronary artery disease, at least for the fixed Hct considered in the study.
Regarding studies in which a wide range of hematocrits is analyzed, covering healthy and non-physiological patients, in [39], Farias et al. conducted a numerical study using CFD simulations to investigate the influence of Hct levels (ranging from 5% to 50%) on blood dynamics. They employed a power-law viscosity model in a simplified carotid artery geometry, and their results were compared against a fixed Newtonian viscosity model representative of a healthy subject. The study revealed notable differences in flow dynamics across various Hct levels, evaluated in terms of vortex shedding frequency, velocity range, and flow bifurcation between the carotid outlet branches. Furthermore, the findings highlighted similarities between the non-Newtonian simulations and the Newtonian case within the healthy Hct range, whereas significant differences were observed at Hct levels outside this range. Similarly, in [40], Lynch et al. investigated the effects of non-Newtonian blood viscosity on hemodynamic and transport metrics in patient-specific arterial and venous geometries. Their findings revealed that non-Newtonian blood behavior significantly alters flow dynamics, particularly in regions of low shear rate and recirculation, such as aneurysms and large veins. Compared to Newtonian models, non-Newtonian models exhibited lower in-plane velocity and vorticity but higher WSS. Transport analyses also demonstrated differences in scalar concentration and particle residence time, highlighting the importance of accurate rheological modeling to better understand cardiovascular mass transport and disease processes, including thrombosis and atherosclerosis.
Previous studies have also explored modeling and simulating blood flow using more complex rheological models to account for local phenomena, such as the Fahraeus effect [41]. These approaches include considering a two-layered fluid composed of a biphasic layer (plasma and cells) and a plasmatic layer (only plasma) [42], or employing alternative blood models such as the Herschel–Bulkley fluid [43].
Further investigations have also focused on evaluating the impact of blood rheology in Fluid-Structure Interaction (FSI) simulations of the great vessels. For instance, Qiao et al. [44] investigated the performance of the Quemada, Casson, Carreau, and Carreau-Yasuda models in simulating blood flow in healthy aortas using FSI methods. Their findings showed that instantaneous low shear strain rates (SSR $<$ 100 s-1), where the non-Newtonian blood assumption is valid, constitute a significant proportion of the cardiac cycle and aortic wall, potentially influencing blood viscosity considerably. Furthermore, they demonstrated that high-viscosity regions, consistently observed in the descending aorta, emphasize the necessity of non-Newtonian models for accurate aortic hemodynamics simulations.
There have also been interest in incorporating non-Newtonian viscosity models to estimate hemodynamic parameters from velocity measurements obtained using 4D Flow MRI. For instance, in [45], Vaclavu et al. evaluated the hemodynamic within the circle of Willis in children and adults with Sickle Cell Disease (SCD) using 4D Flow MR imaging and a non-Newtonian rheology model. Their results showed that SCD patients had lower Hct and viscosity, and higher velocity, flow and lumen area, with lower Endotelial Shear Stress (ESS) compared to healthy controls. When comparing only adults, they found that SCD patients shared similar velocity ranges compared to healthy controls, but lower ESS.
In [46], Riva et al. investigated the effect of three viscosity models on the quantification of patient-specific viscous energy loss ($E_{L}$) from 4D Flow MRI within the left ventricle (LV): a fixed Newtonian viscosity of 3.7 cP, Newtonian viscosity adjusted to patient-specific Hct, and the Quemada model adjusted to patient-specific Hct. Their results indicated that the calculation of viscosity-dependent parameters, and consequently the characterization of LV blood energetics using 4D Flow MRI, can be influenced by the chosen viscosity model. This effect was particularly pronounced in regions with recirculation or low velocities. Although the study focused solely on LV diastolic dysfunction, the authors noted that their conclusions could also apply to other conditions characterized by increased recirculating blood in the LV, such as ischemic cardiomyopathy, or by non-pulsatile pulmonary blood flow, as observed in post-Fontan procedure cases. Furthermore, the authors emphasized the importance of considering Hct derangements, which are often associated with cardiovascular risk factors such as hypertension, diabetes, and obesity, among others.
In [47], Cheng and colleagues conducted in-silico studies involving geometries resembling Fontan patients, as well as in-vitro 4D Flow MRI phantom experiments, to evaluate the impact of shear-dependent fluid viscosity on in-vitro Fontan circulation. The investigation utilized one idealized and two patient-specific 3D-printed Fontan circuit models, which were connected to a pulsatile pump with parameters tuned to mimic cardiac circulation. The three phantoms were measured twice using 4D Flow MRI to obtain velocity maps: once with a Newtonian fluid and once with a non-Newtonian fluid resembling blood. The results demonstrated that in all three models, the non-Newtonian fluid with shear-dependent viscosity significantly altered flow patterns, power loss, flow distribution, and flow mixing compared to the Newtonian fluid. Moreover, the authors suggested that the observed differences could be associated with clinically impactful outcomes, emphasizing the importance of considering non-Newtonian fluid properties in hemodynamic analyses.
A similar study evaluated blood flow through a Bileaflet Mechanical Heart Valve (BMHV) using FSI simulations validated with in-vitro experimental data obtained through Particle Imaging Velocimetry [48]. The study was conducted within a heart model designed to assess the performance and functionality of prosthetic heart valves. The authors found that variations in Hct significantly influence blood dynamics, particularly shear stress distribution and flow recirculation within the aortic sinus. They also demonstrated that increased Hct amplifies shear-thinning behavior, leading to non-linear changes in shear stress and altered valve leaflet dynamics, including asymmetric opening and closure patterns. These findings highlight the importance of incorporating patient-specific Hct values into computational models to improve predictions of BMHV performance and support personalized medical assessments.
Although there is disparity among the cited authors regarding the necessity of non-Newtonian blood modeling, it is important to note that most of their work has primarily focused on the healthy Hct range. Consequently, their findings cannot be generalized to cases involving abnormal blood composition. However, studies evaluating a wider range of Hct values consistently conclude that rheology plays a significant role, regardless of the specific non-Newtonian model used.
@@
Despite the valuable insights provided by the aforementioned studies, several open questions remain. For instance, the impact of neglecting shear-thinning blood behavior when estimating hemodynamic parameters from 4D Flow MRI velocity measurements for the assessment of blood hemodynamics remains unclear, particularly in the aorta, where non-Newtonian blood behavior is often overlooked by the MRI community. Furthermore, none of these studies have assessed whether the influence of shear-thinning blood behavior remains observable under common MR imaging conditions, such as partial volume effects, flow artifacts, and noise, when using 4D Flow MRI velocity measurements.
To address these open questions, in this investigation, we: (1) evaluated whether patient-specific Newtonian and non-Newtonian viscosities yield equivalent results in terms of WSS, $\dot{E}_{L}$ (rate of viscous energy loss dissipation), and OSI across different Hct levels when estimated from 4D Flow MR velocity measurements; (2) quantified the differences between these parameters when estimated using patient-specific non-Newtonian and state-of-the-art Newtonian viscosities commonly reported in the literature; and (3) showed that if blood exhibits shear-thinning behavior, this effect remains observable in 4D Flow MRI data, even in the presence of imaging artifacts under common imaging conditions.
To achieve this, we conducted numerical and in-vivo experiments to estimate hemodynamic parameters from 4D Flow MR images using a power-law non-Newtonian viscosity model and Newtonian viscosities for blood rheology, comparing WSS, $\dot{E}_{L}$, and OSI estimations. This study integrates the 3D Cauchy momentum equations (often loosely referred to as the Navier-Stokes equations when accounting for non-Newtonian fluids) with 0D surrogate models for downstream flow and non-linear stress tensors to simulate shear-thinning blood behavior. CFD simulations were used to generate synthetic 4D Flow MR images, which, together with in-vivo 4D Flow MRI data from five patients with Hypertrophic Cardiomyopathy, were analyzed for direct estimation of hemodynamic parameters using image processing methodologies. In both cases, results obtained using power-law viscosities adjusted to each patient’s Hct (measured at physiological temperatures) were compared with those derived from Hct-fitted and state-of-the-art Newtonian viscosities commonly reported in the CMR literature.
We believe that this investigation could significantly impact the current estimation of hemodynamic parameters from MR images in clinical practice, as incorporating more realistic viscosity models may enhance the assessment of cardiovascular diseases. This is particularly relevant for patients with conditions affecting their Hct (such as those mentioned in earlier), which can further influence their blood dynamics. Finally, this study paves the way for integrating patient-specific, Hct-based non-Newtonian viscosities into future MRI-based diagnostic workflows.
@@ 2.1 Mathematical modeling
Let $\Omega\in\mathbb{R}^{3}$ be the three-dimensional computational domain, with its two-dimensional boundaries decomposed into disjoint subsets such that $\partial\Omega=\Gamma_{\text{in}}\cup_{k=1}^{l}\Gamma_{\text{out}}^{k}\cup\Gamma_{\text{wall}}$, where $\Gamma_{\text{in}}$ represents the inlet and $\Gamma_{\text{out}}$ refers to the outlets, and $\Gamma_{\text{wall}}$ denotes the artery walls. The incompressible Cauchy momentum equations relate the velocity vector field ($\boldsymbol{u}$) and the pressure ($p$) field as follows:
@@
where $k=1,\ldots,l$ is the outlet index, $\tau$ denotes the shear stress tensor of the fluid, and $\rho$ is the fluid density. The physiological value $\rho=1.06$ gr///cm3 was adopted in this study. Modeling choices for $\tau$ are discussed in the following section. Additionally, $g(x,t)$ is a is a spatially and temporally dependent boundary condition used to define the pulsatile behavior of blood flow.
To model the downstream fluid, we coupled the 3D Cauchy momentum equations with reduced 0D three-parameters Windkessel models [49] (see Figure 1), as commonly done in the literature [50, 51, 52, 53]. This approach reduces to solving an Ordinary Differential Equation (ODE) for each outlet domain:
$p^{k}_{\text{wk}}=R_{p}Q^{k}(\boldsymbol{u})+p^{k}_{d},\quad k=1,\ldots,l,$ (2.2a)
$C^{k}\frac{dp^{k}_{d}}{dt}+\frac{p^{k}_{d}}{R^{k}_{d}}=Q^{k}(\boldsymbol{u}),\quad k=1,\ldots,l.$ (2.2b)
@@
Figure 1: Simulation workflow for generating 4D Flow MR images. First, the aortic model serves as input to the CFD solver, where the Cauchy momentum equations with nonlinear viscosity are solved. Subsequently, $k$-space locations and sequence timings are generated using parameters from Table 3, and the signal equation in (2.4) is evaluated. Trapezoidal MRI gradients are considered for the generation of sequence timings.